Factor completely over the integers. Check the answer using multiplication.
(y+3)^2-4y^2

Found 2 solutions by Fombitz, ikleyn:
Answer by Fombitz(32388)
Answer by ikleyn(52800)

.
(y+3)^2 - 4y^2 = (y+3)^2 - (2y)^2 = 


    factor as the difference of squares a^2 - b^2  = (a+b)*(a-b) 


= (y+3 + 2y)*(y+3-2y) = (3y+3)*(-y+3) = -3*(y+1)*(y-3).